An AC source powers a circuit containing a resisto
Answer #1

capacitance, C=1F,


resistance, R=1 ohm


inductance L=1H


and w=0.5Hz, and wo=1Hz

impedance Z=sqrt(R^2+(XL-XC)^2)


Z=sqrt(R^2+(w*L-1/w*c)^2)


Z=sqrt(1^2+(0.5*1-1/(0.5*1))^2)


Z=1.803 ohms